An urn contains 3 white balls and 5 red balls. Two balls are drawn at random. If the first is replaced before the second is drawn, what is the probability that the first ball is white and the second is red?
Answer by drk(1908)

P(white then red) = 3/8*5/8 = 15/64.
Since we are replacing, the denominator stays the same.